OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 3y = 4x + 7 5y = 4x + 25 A. (2, 9) B. (2, 5) C. (5, 5) D. (5, 9)

OpenStudy (anonymous):

diddnt get anything for this one

OpenStudy (anonymous):

a

OpenStudy (anonymous):

elimination is the easier method here. put all x's and y's on the left side of the eq so you get 3y-4x=7 5y-4x=25 now multiply either eq by -1 so -3y+4x=-7 5y-4x=25 now add the eqs and you get 2y+0x=18 y=9 now plug y=9 into either eq and see what you get for x